# Multi-task neural diffusion processes for uncertainty-quantified wind power prediction

### v1

@@ 6.4.2 Model comparison
Fig. 13(a) compares empirical and nominal coverage for the same models. The five-dimensional MT-NDP demonstrates the best calibration overall, with coverage curves lying closest to the diagonal as context size increases. By contrast, the one-dimensional ST-NDP shows little benefit from context, underscoring its limited ability to capture multi-feature relationships.
Finally, Fig. 13(b) summarises these findings using the CE. The five-dimensional MT-NDP again exhibits the strongest performance, achieving lower CE values than both the one-dimensional and five-dimensional ST-NDPs, particularly at higher context sizes. These results confirm that multi-task extension improves both point prediction accuracy and uncertainty calibration in a challenging case of turbine $6$.
@@
Table 2: Out-of-sample performance of single-task neural diffusion process (ST-NDP) and multi-task neural diffusion process (MT-NDP) models on turbine $6$, averaged over $30$ test functions ($100$ points each). Metrics reported: mean absolute error (MAE, kW), root mean squared error (RMSE, kW), and coverage error (CE, %).
Model # of features Context MAE (kW) RMSE (kW) CE (%)
ST-NDP 11 0 $66.539$ $90.491$ $13.052$
$25$ $56.641$ $77.766$ $8.779$
$50$ $53.283$ $73.429$ $6.865$
ST-NDP 55 0 $55.169$ $75.441$ $11.225$
$25$ $51.864$ $71.306$ $9.058$
$50$ $49.676$ $68.465$ $7.376$
MT-NDP 55 0 $55.291$ $75.974$ $11.375$
$25$ $49.183$ $67.805$ $6.986$
$50$ $47.268$ $65.636$ $5.691$
@@ 7 Discussion
7.1 Key findings
Our results show three consistent patterns. First, in the one-dimensional setting where wind speed is the sole input feature, NDPs match GPs in point accuracy while providing markedly better uncertainty calibration. As seen in Fig. 7, MAE and RMSE distributions are similar for the two models, but the NDP’s empirical coverage lies substantially closer to the $y=x$ line than the GP’s, which is underconfident across coverage levels. This aligns with prior evidence that GP Gaussianity can yield miscalibrated intervals on SCADA data.
Second, increasing the input dimensionality from one dimension to five dimensions, i.e., by adding sine/cosine wind direction, nacelle temperature, and transformer temperature, improves point accuracy for NDPs, and the benefit grows with more context. Fig. 8 shows lower MAE/RMSE for the five-dimensional NDP model, with clear gains as context increases from 0 to $50$ points. Calibration also improves with context (Fig. 10), narrowing the gap to the one-dimensional NDP model at $50$ points. These trends suggest that richer covariates and larger context sets are complementary: the former increase explanatory power; the latter help the model specialise to the test turbine.
Third, extending NDPs to the multi–task setting (MT–NDP) yields the strongest performance. The five-dimensional MT–NDP outperforms the one-dimensional and five-dimensional ST-NDPs in both accuracy and calibration as the context grows (Fig. 12, 13; Table 2), demonstrating an effective adaptation of a few shots to the behaviour of the out-of-distribution turbine.
7.2 Limitations
Our experimental design has three main limitations. (i) Context budget: we evaluated context sizes up to $50$ points. Given the consistent gains for the five-dimensional NDP models, larger context sets could further improve calibration and may even overturn the slight coverage advantage occasionally observed for the one-dimensional model at $50$ points. (ii) Encoder simplicity: the task encoder is an MLP chosen for input–size agnosticity and permutation equivariance via averaging; it does not guarantee invariance across feature orderings as attention-based designs do. (iii) Data homogeneity: the Kelmarsh turbines are relatively homogeneous in the one-dimensional regime, which limits how much benefit context can bring and how strongly multi–task methods are stress-tested. Additionally, rare operating regions can produce occasional loss outliers, suggesting that stratified sampling could help.
7.3 Future work
Two immediate extensions are computational rather than conceptual: (i) increasing the context budget for higher–dimensional models and (ii) expanding the feature set beyond five SCADA variables by leveraging GPUs for training and sampling. Beyond these, we identify three methodological directions: (a) attention-based or bi-dimensional attention encoders to enforce invariance and capture richer task structure; (b) time-aware NDP variants that account for temporal dependence within functions; and (c) broader multi–farm evaluations—including onshore and offshore sites with diverse turbine types—to assess robustness and transferability.
7.4 Implications
Taken together, the results indicate that NDPs provide calibrated, uncertainty-aware predictions, scale effectively to richer inputs, and—when equipped with a task encoder—deliver few–shot adaptation even to turbines that deviate from the fleet average. In operational settings, this combination of accuracy and calibration matters: sharper yet trustworthy intervals support dispatch, bidding, and maintenance decisions under uncertainty. Focusing on genuinely distinct turbines, such as turbine $6$ in our study, offers the clearest evidence of these advantages in practice.
@@ 8 Conclusion
This paper has presented the first empirical evaluation of NDPs and their multi-task extension for wind power prediction from SCADA data. We have shown that NDPs match the predictive accuracy of GPs while substantially outperforming them in terms of calibrated uncertainty estimates. Unlike GPs, NDPs scale efficiently to higher-dimensional input spaces and remain tractable on large datasets, making them suitable for real-world deployment.
A key strength of NDPs is their ability to generalise to unseen turbines with only a handful of context points, enabling few-shot inference in settings where historical data are scarce. Extending the framework to the multi-task case, we introduced an MT-NDP architecture with a task encoder and demonstrated that it can exploit cross-turbine correlations to further improve accuracy and uncertainty calibration. These benefits are particularly evident on wind turbines that deviate most strongly from the fleet average, where single-task models often struggle or cannot be applied effectively.
From an operational perspective, these findings are highly relevant for wind farm monitoring and forecasting. Reliable uncertainty quantification supports decision-making for grid integration, market participation, and maintenance scheduling. By improving predictive robustness and reducing reliance on large task-specific datasets, NDP-based models could help reduce downtime and operating costs in modern wind farms.
Future work will focus on scaling MT-NDPs to the full set of SCADA variables, incorporating temporal dependencies to better capture dynamic turbine behaviour, and testing on more heterogeneous wind farms, including offshore installations. Enhancing the task encoder with attention mechanisms also offers a promising direction for capturing richer cross-turbine structure.

### v2

@@ 5.2.2 Model comparison
Fig. 8(a) compares empirical and nominal coverage for the same models. The five-dimensional MT-NDP demonstrates the best calibration overall, with coverage curves lying closest to the diagonal as context size increases. By contrast, the one-dimensional ST-NDP shows little benefit from context, underscoring its limited ability to capture multi-feature relationships. Finally, Fig. 8(b) summarises these findings using the CE. The five-dimensional MT-NDP again exhibits the strongest performance, achieving lower CE values than both the one-dimensional and five-dimensional ST-NDPs, particularly at higher context sizes. These results confirm that multi-task extension improves both point prediction accuracy and uncertainty calibration in a challenging case of turbine $6$.
@@
Table 1: Out-of-sample performance of single-task neural diffusion process (ST-NDP) and multi-task neural diffusion process (MT-NDP) models on turbine $6$, averaged over $30$ test functions ($100$ points each). Metrics reported: mean absolute error (MAE, kW), root mean squared error (RMSE, kW), and coverage error (CE, %). Bold values indicate the best overall performance across all models and context sizes. 1D: one-dimensional, 5D: five-dimensional.
Context = 0 Context = $25$ Context = $50$
Model MAE RMSE CE MAE RMSE CE MAE RMSE CE
1D ST-NDP $66.539$ $90.491$ $13.052$ $56.641$ $77.766$ $8.779$ $53.283$ $73.429$ $6.865$
5D ST-NDP $55.169$ $75.441$ $11.225$ $51.864$ $71.306$ $9.058$ $49.676$ $68.465$ $7.376$
5D MT-NDP $55.291$ $75.974$ $11.375$ $49.183$ $67.805$ $6.986$ $\mathbf{47.268}$ $\mathbf{65.636}$ $\mathbf{5.691}$
@@ 6 Discussion
We examine the role of task conditioning in diffusion-based function models and its effect on generalisation and uncertainty estimation in multi-task regression. The results show that conditioning on limited context enables adaptation from a shared prior while preserving uncertainty calibration. Introducing a task encoder improves few-shot adaptation by enabling the model to distinguish tasks that deviate from the training distribution, while leaving the underlying diffusion process unchanged. This separation between diffusion-based function modelling and task encoding allows task-specific information to be incorporated without modifying the underlying generative process, preserving the probabilistic structure of the model. Several limitations point to directions for future research. The task encoder is deliberately simple and does not enforce full invariance; larger context budgets were not explored, and the experimental setting exhibits limited heterogeneity in some regimes. Addressing these aspects would clarify how architectural choices influence diffusion-based learning in multi-task settings.
@@ 7 Conclusion
We introduced MT-NDPs as a principled extension of diffusion-based function models that enable task-conditioned probabilistic regression. By augmenting NDPs with a lightweight task encoder, the proposed framework supports few-shot adaptation across related tasks while preserving calibrated uncertainty estimates. Empirical results on wind turbine SCADA data show that MT-NDPs consistently outperform single-task counterparts as context increases, particularly for tasks that deviate from the training distribution. These findings demonstrate that task structure can be incorporated into diffusion-based regression without altering the core diffusion mechanism, highlighting MT-NDPs as a flexible and scalable approach for multi-task learning in function space.
Appendix A Gaussian Processes (GPs)
A GP $f:\mathbb{R}^{D}\rightarrow\mathbb{R}$ is a stochastic process such that, for any finite collection of inputs $X=\{x_{i}\}_{i=1}^{N}\subset\mathbb{R}^{D}$, the vector of function evaluations $f(X)=[f(x_{1}),\ldots,f(x_{N})]^{T}$ follows a multivariate normal distribution [Rasmussen2006GaussianLearning]. GPs satisfy the Kolmogorov extension theorem, which guarantees that all finite-dimensional marginals are mutually consistent under permutation and marginalisation [kolmogorov1950]. This provides a principled Bayesian framework for modelling distributions over functions, enabling exact inference with predictive uncertainty. In our experiments presented in Section 5, we use GP regression as a benchmark model, so we briefly outline its formulation here.
$f(X)\sim\mathcal{N}(\mu(X),K(X,X)),$ (15)
$E(y^{*}|y)$ $=\mu(X^{*})+K(X^{*},X)A^{-1}(y-\mu(X)),$ (16)
$V(y^{*}|y)$ $=K(X^{*},X^{*})-K(X^{*},X)A^{-1}K(X,X^{*}).$ (17)
For our experiments, we use the RBF kernel:
Appendix B Diffusion probabilistic models (DPMs)
DPMs were introduced as a class of generative models inspired by non-equilibrium thermodynamics [sohldickstein2015deepunsupervisedlearningusing]. The central idea is to define a forward process, in which the structure of the data distribution is gradually destroyed by the iterative addition of Gaussian noise, and a corresponding reverse process that learns to invert this corruption.
The de-noising diffusion probabilistic model (DDPM) [ho2020denoisingdiffusionprobabilisticmodels] simplified and stabilised training and has become the standard formulation. Following [ho2020denoisingdiffusionprobabilisticmodels, dutordoir2023neural], the forward process begins from the input data distribution $q(s_{0})$, where $s_{0}\in\mathbb{R}^{D}$ denotes a $D$-dimensional data vector. It defines a fixed Markov chain over the sequence $s_{0:T}=\{s_{0},\ldots,s_{T}\}$ with density
$q(s_{0:T})=q(s_{0})\prod_{t=1}^{T}q(s_{t}\mid s_{t-1})$ (19)
and conditional distributions
where $\{\beta_{t}\in(0,1)\}_{t=1}^{T}$ is a variance schedule controlling the magnitude of injected noise and $I_{D}$ is the $D\times D$ identity matrix. After $T$ steps, the distribution approaches Gaussian noise, $q(s_{T})\approx\mathcal{N}(0,I)$. Early implementations used a linear variance schedule, while later work demonstrated that a cosine schedule improves performance by distributing noise more evenly across diffusion steps [nichol2021improveddenoisingdiffusionprobabilistic]. A variance schedule defines the cumulative noise parameter
$\bar{\alpha}_{t}=\prod_{j=1}^{t}(1-\beta_{j}),$ (21)
The reverse process is intractable in closed form, but can be approximated by a NN that predicts the injected noise. The mean of the reverse transition is parameterised as
where $\alpha_{t}=1-\beta_{t}$, $\bar{\alpha}_{t}=\prod_{j=1}^{t}\alpha_{j}$, and $\epsilon_{\theta}:\mathbb{R}^{D}\times\{1,\dots,T\}\to\mathbb{R}^{D}$ is an NN that predicts the injected noise vector. Training is carried out by minimising the score-matching loss
This framework enabled diffusion models to achieve state-of-the-art generative performance, rivalling generative adversarial networks (GANs) and autoregressive models. Building on DDPMs, Repaint [lugmayr2022repaintinpaintingusingdenoising] adapted the approach to image inpainting by conditioning on context points during the reverse process, without altering the pre-trained DDPM. This method achieved high-quality and diverse reconstructions under challenging conditions, outperforming autoregressive and GAN-based baselines.
In parallel, NPs were proposed as a stochastic process model combining NNs with an encoder–decoder architecture [garnelo2018neuralprocesses]. NPs learn to map a set of context points into a latent task representation, which is then used by a decoder to generate predictions at target inputs. This provides scalability and the ability to adapt to new tasks with only a few observations, but standard NPs often underfit and produce poorly calibrated uncertainty [kim2019attentiveneuralprocesses]. Latent and attentive variants improve correlations between targets, but still fall short of fully consistent Bayesian behaviour [dutordoir2023neural].
NDPs unify these two developments, drawing on diffusion-based generative modelling to overcome the Gaussian assumptions of NPs while retaining their scalability and adaptability. NDPs extend diffusion models from distributions over data to distributions over functions [dutordoir2023neural], thereby enabling uncertainty-aware regression.
Appendix C Dataset description
An operational status and events file was used to filter the SCADA data and ensure consistent modelling of normal turbine behaviour. These logs provide detailed information on turbine operating states, including technical failures as well as operational or environmental standbys and warnings. To enable accurate modelling, all data associated with out-of-control conditions were removed so that model training was based solely on stable operating periods. Specifically, records corresponding to standbys, warnings, and operational stops were excluded. In addition, data from the week preceding each forced outage was discarded to reduce the likelihood of capturing deteriorating states. The resulting data elimination process is also described and illustrated in [Fiocchi_2025].
Figure 9: SCADA data for the six turbines from the Kelmarsh wind farm [Plumley2022] after removing standbys and warnings using the operational status and events file. (a) Wind power curves showing mean power output (kW) by mean wind speed (m/s). (b) Joint distribution of mean wind speed (m/s) and mean wind direction (°) for turbine 1 (left), and mean power output (kW) by mean wind speed (m/s) and wind direction (°) for turbine 1 (right).
Figure 10: SCADA data for the six turbines from the Kelmarsh wind farm [Plumley2022] after removing standbys and warnings using the operational status and events file. (a) Mean power output (kW) by mean nacelle temperature (°C). (b) Mean power output (kW) by mean transformer temperature (°C).
Appendix D 1D NDP against 5D NDP
Figure 11: Distributions of mean absolute error (MAE, left) and root mean square error (RMSE, right) for the one-dimensional (1D) and five-dimensional (5D) neural diffusion process (NDP) models, evaluated with 0, $25$, and $50$ context points. Results are aggregated over $30$ test functions of $100$ points each from turbine $1$. ST: single-task.
Figure 12: Observed vs. theoretical coverage probabilities for the one-dimensional (1D) and five-dimensional (5D) neural diffusion process (NDP) models with 0, $25$, and $50$ context points. Results are averaged over $30$ test functions of $100$ points each from turbine $1$. ST: single-task.
Table 2: Out-of-sample performance of Gaussian process (GP) and single-task neural diffusion process (ST-NDP) models on turbine $1$, averaged over $30$ test functions ($100$ points each). Metrics reported: mean absolute error (MAE, kW), root mean squared error (RMSE, kW), and coverage error (CE, %). 1D: one-dimensional, 3D: three-dimensional, 5D: five-dimensional.
Context = 0 Context = $25$ Context = $50$
Model MAE RMSE CE MAE RMSE CE MAE RMSE CE
1D GP $51.636$ $75.596$ $18.075$
1D ST-NDP $52.603$ $75.139$ $4.584$ $51.273$ $74.568$ $4.181$ $50.545$ $74.137$ $3.938$
3D ST-NDP $60.967$ $88.096$ $9.731$ $52.338$ $74.312$ $6.750$ $50.855$ $73.537$ $5.369$
5D ST-NDP $49.429$ $67.685$ $8.351$ $46.962$ $66.007$ $6.029$ $45.791$ $65.301$ $4.816$